OpenStudy (goformit100):

An edge of a variable cube is increasing at the rate of 3 cm/s. How fast is the volume of the cube increasing when the edge is 10 cm long?

OpenStudy (goformit100):

@terenzreignz

terenzreignz (terenzreignz):

okay... if we let x be the length of one edge of the cube, we get that the volume of the cube is equal to... \[\huge V = x^3\] Now... we're looking for \[\huge \left.\frac{dV}{dt}\right.\]

terenzreignz (terenzreignz):

Differentiating seems simplest... \[\huge \frac{dV}{dt}=\frac{dV}{dx}\times \frac{dx}{dt}\] And we do have a value for the dx/dt... it's just 3. So once this is done, plug in x = 10, and you're all set.

OpenStudy (goformit100):

Thankyou sir